OpenStudy (anonymous):

Which mixed number falls between four and three-eighths and five and five-twelfths? four and one-fourth five and two-thirds four and one-half five and five-sixths

OpenStudy (anonymous):

Add the two numbers and then divide by two \[(4+3/8+5+5/12)/2\]

OpenStudy (anonymous):

\[4 + \frac{3}{8} < x < 5 + \frac{5}{12}\]

OpenStudy (anonymous):

\[4+\frac{1}{2}\]

OpenStudy (anonymous):

And The Answer Is?

OpenStudy (anonymous):

Get a calculator and punch in my answer, I spelled it out for you.

OpenStudy (anonymous):

We know 4 and one-fourth is smaller than five (and five-twefths). Is it larger than four and three-eighths? This is the same thing as asking is one-fourth larger than three-eights?\[\frac{1}{4}=\frac{2}{8}<\frac{3}{8}\]Thus the answer is not four and one-fourth. Going through the other three we find the answer is four and one-half.

OpenStudy (anonymous):

\[\left\{4+\frac{3}{8},5+\frac{5}{12}\right\}=\left\{\frac{35}{8},\frac{65}{12}\right\}=\{4.375,5.41667\} \]Using there above one can virtually make a mental comparison with the four answer offerings.